Calculate the photon energy of a radio transmission with a frequency of 93.9 MHz.
1
Expert's answer
2019-05-02T07:31:33-0400

"E=h\nu"

E - photon energy, J

h - Planck's constant, "h \approx 6.626 \cdot 10^{-34} \; J \cdot s"

"\nu" - frequency, Hz

"E=6.626 \cdot 10^{-34} \cdot 93.9 \cdot 10^6 \approx 6.22 \cdot 10^{-26} \; J"


Answer: "6.22 \cdot 10^{-26} \; J"